# Test Time Learning for Time Series Forecasting

### v2

@@ 5.1 Experimental Setup with Enhanced Architectures
To assess the impact of enhancing the model architecture, we conducted experiments by adding hidden layer architectures before the sequence modeling block in each of the four TTT blocks. The goal was to improve performance by enriching feature extraction through local temporal context. As shown in Figure 3 in Appendix B.
@@
ModernTCN showed competitive results across multiple datasets (see Appendix D), such as ETTh2, where it achieved an MSE of 0.285 at horizon 96, outperforming Conv 3 and Conv 5. However, as with other deep convolutional layers, ModernTCN’s increased complexity also led to slower training times compared to simpler setups like Conv 3 and it failed to match Conv Stack 5’s performance.
@@ 5.3 Results and Statistical Comparisons for Proposed Architectures
The proposed architectures—TTT Linear, TTT MLP, Conv Stack 3, Conv Stack 5, Conv 3, Conv 5, Inception, and TTT with ModernTCN—exhibit varying performance across prediction horizons. TTT Linear performs well at shorter horizons (MSE 0.268, MAE 0.298 at horizon 96) but declines at longer horizons (MSE 0.357 at horizon 336). TTT MLP follows a similar trend with slightly worse performance. Conv 3 and Conv 5 outperform Linear and MLP at shorter horizons (MSE 0.269, MAE 0.297 at horizon 96) but lag behind Conv Stack models at longer horizons. TTT with ModernTCN shows promising results at shorter horizons, such as MSE 0.389, MAE 0.402 on ETTh1, and MSE 0.285, MAE 0.340 on ETTh2 at horizon 96. Although results for Traffic and Electricity datasets are pending, preliminary findings indicate TTT with ModernTCN is competitive, particularly for short-term dependencies (see Table 7 in Appendix D). Conv Stack 5 performs best at shorter horizons (MSE 0.261, MAE 0.289 at horizon 96). Inception provides stable performance across horizons, closely following Conv Stack 3 (MSE 0.361 at horizon 336). At horizon 720, Conv 5 shows a marginal improvement over Conv 3, with an MSE of 0.400 compared to 0.406. The Conv Stack 5 architecture demonstrates the best overall performance among all convolutional models.
@@ 5.5 Evaluation
  • Impact of Hidden Layer Architectures: While stacked convolutional architectures, such as Conv Stack 3 and Conv Stack 5, provided incremental improvements, simpler architectures like Conv 3 and Conv 5 also delivered competitive performance. Conv Stack 5 showed a reduction in MSE compared to TimeMachine, at horizon 96, where it achieved an MSE of 0.261 versus TimeMachine’s 0.262. The ModernTCN architecture showed promising results, especially in capturing short-term dependencies effectively.
@@ 6 Conclusion and Future Work
In this work, we improved the state-of-the-art (SOTA) model for time series forecasting by replacing the Mamba modules in the original TimeMachine model with Test-Time Training (TTT) modules, which leverage linear RNNs to capture long-range dependencies. Extensive experiments demonstrated that the TTT architectures—MLP and Linear—performed well, with MLP slightly outperforming Linear. Exploring alternative architectures, particularly Conv Stack 5 and ModernTCN, significantly improved performance at longer prediction horizons, with ModernTCN showing notable efficiency in capturing short-term dependencies. The most significant gains came from increasing sequence and prediction lengths, where our TTT models consistently matched or outperformed the SOTA model, particularly on larger datasets like Electricity, Traffic, and Weather, emphasizing the model’s strength in handling long-range dependencies. While convolutional stacks and ModernTCN showed promise, further improvements could be achieved by refining hidden layer configurations and exploring architectural diversity. Overall, this work demonstrates the potential of TTT modules in long-term forecasting, especially when combined with robust convolutional architectures and applied to larger datasets and longer horizons.
Appendix A Appendix A: TTT vs Mamba
@@
A.1 Mamba: Gated Linear RNN via State Space Models (SSMs)
@@
A.2 TTT: Test-Time Training with Dynamic Hidden States
@@
A.3 Comparison of Complexity and Adaptability
@@
Appendix B Appendix B: Model Components
B.1 TTT Block and Proposed Architectures
@@ B.3 Normalization
Appendix C Appendix C: Analysis on Increased Prediction and Sequence Length
C.1 Effect of Sequence Length

### v3

@@ 5.1 Experimental Setup with Enhanced Architectures
To assess the impact of enhancing the model architecture, we conducted experiments by adding hidden layer architectures before the sequence modeling block in each of the four TTT blocks. The goal was to improve performance by enriching feature extraction through local temporal context. As shown in Figure 3 in Appendix D.
@@
ModernTCN showed competitive results across multiple datasets (see Appendix I), such as ETTh2, where it achieved an MSE of 0.285 at horizon 96, outperforming Conv 3 and Conv 5. However, as with other deep convolutional layers, ModernTCN’s increased complexity also led to slower training times compared to simpler setups like Conv 3 and it failed to match Conv Stack 5’s performance.
@@ 5.3 Results and Statistical Comparisons for Proposed Architectures
The proposed architectures—TTT Linear, TTT MLP, Conv Stack 3, Conv Stack 5, Conv 3, Conv 5, Inception, and TTT with ModernTCN—exhibit varying performance across prediction horizons. TTT Linear performs well at shorter horizons (MSE 0.268, MAE 0.298 at horizon 96) but declines at longer horizons (MSE 0.357 at horizon 336). TTT MLP follows a similar trend with slightly worse performance. Conv 3 and Conv 5 outperform Linear and MLP at shorter horizons (MSE 0.269, MAE 0.297 at horizon 96) but lag behind Conv Stack models at longer horizons. TTT with ModernTCN shows promising results at shorter horizons, such as MSE 0.389, MAE 0.402 on ETTh1, and MSE 0.285, MAE 0.340 on ETTh2 at horizon 96. Although results for Traffic and Electricity datasets are pending, preliminary findings indicate TTT with ModernTCN is competitive, particularly for short-term dependencies (see Table 7 in Appendix I). Conv Stack 5 performs best at shorter horizons (MSE 0.261, MAE 0.289 at horizon 96). Inception provides stable performance across horizons, closely following Conv Stack 3 (MSE 0.361 at horizon 336). At horizon 720, Conv 5 shows a marginal improvement over Conv 3, with an MSE of 0.400 compared to 0.406. The Conv Stack 5 architecture demonstrates the best overall performance among all convolutional models.
@@ 5.5 Evaluation
  • Impact of Hidden Layer Architectures: While stacked convolutional architectures, such as Conv Stack 3 and Conv Stack 5, provided incremental improvements, simpler architectures like Conv 3 and Conv 5 also delivered competitive performance. Conv Stack 5 showed a reduction in MSE compared to TimeMachine, at horizon 96, where it achieved an MSE of 0.261 versus TimeMachine’s 0.262. ModernTCN failed to meet the performance of simpler architectures.
@@ 6 Conclusion and Future Work
In this work, we improved the state-of-the-art (SOTA) model for time series forecasting by replacing the Mamba modules in the original TimeMachine model with Test-Time Training (TTT) modules, which leverage linear RNNs to capture long-range dependencies. Extensive experiments demonstrated that the TTT architectures—MLP and Linear—performed well, with MLP slightly outperforming Linear. Exploring alternative architectures, particularly Conv Stack 5 and ModernTCN, significantly improved performance at longer prediction horizons, with ModernTCN showing notable efficiency in capturing short-term dependencies. The most significant gains came from increasing sequence and prediction lengths, where our TTT models consistently matched or outperformed the SOTA model, particularly on larger datasets like Electricity, Traffic, and Weather, emphasizing the model’s strength in handling long-range dependencies. While convolutional stacks and ModernTCN showed promise, further improvements could be achieved by refining hidden layer configurations and exploring architectural diversity. We included some potential real world applications of the TTT module in Appendix C along with why we believe it’s best suited for LTSF. Overall, this work demonstrates the potential of TTT modules in long-term forecasting, especially when combined with robust convolutional architectures and applied to larger datasets and longer horizons.
Appendix A Theory and Motivation of Test-Time Training
A.1 Motivation of TTT on Non-Stationary Data
Time series forecasting often faces challenges arising from non-stationary data, where the underlying statistical properties of the data evolve over time. Traditional models struggle with such scenarios, as they are typically trained on static distributions and are not inherently equipped to handle distribution shifts at inference time. Test-Time Training (TTT) has gained attention as a robust paradigm to mitigate this issue, enabling models to adapt dynamically during inference by leveraging self-supervised learning tasks. For example, the work on self-adaptive forecasting introduced by Google in [3] demonstrates how incorporating adaptive backcasting mechanisms allows models to adjust their predictions to evolving patterns in the data, improving accuracy and robustness under non-stationary conditions. Similarly, FrAug [5] explores data augmentation in the frequency domain to bolster model performance in distributionally diverse settings. While not explicitly a TTT method, FrAug’s augmentation principles align with TTT’s objectives by enhancing model resilience to dynamic changes in time series characteristics. These studies collectively highlight the potential of adaptive methods like TTT to address the unique challenges posed by non-stationary time series data, making them well-suited for applications where robustness and flexibility are paramount.
A.2 Theoretical Basis for TTT’s Adaptability Without Catastrophic Forgetting
TTT avoids catastrophic forgetting by performing online self-supervised learning during inference. The adaptation occurs for each test sample independently, ensuring that the original parameters of the model remain largely intact. The authors that originally proposed TTT provided most of the following mathematical theory in [33] where you can find more detailed explanations.
Mathematical Framework:
Let:
  • $x\in\mathcal{X}$ be the input.
  • $y\in\mathcal{Y}$ be the corresponding label.
  • $f_{\theta}:\mathcal{X}\to\mathcal{Y}$ be the model parameterized by $\theta$.
  • $\mathcal{L}_{\text{main}}(\theta;x,y)$ be the primary task loss.
  • $\mathcal{L}_{\text{self}}(\theta;x)$ be the self-supervised task loss.
At test time, TTT minimizes $\mathcal{L}_{\text{self}}$ for each input $x$, updating the model parameters as:
$\theta^{\prime}=\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{self}}(\theta;x),$
where $\eta>0$ is the learning rate.
Adaptability Without Forgetting:
  • Adaptation is performed on $\mathcal{L}_{\text{self}}$, which does not require labels or the main task’s gradients.
  • Since $\theta^{\prime}$ is computed independently for each test sample, no accumulated parameter updates overwrite prior knowledge.
  • Theoretical support: The optimization of $\mathcal{L}_{\text{self}}$ ensures that changes in parameters $\theta$ are local and transient, i.e., specific to each test sample.
Proof of No Forgetting:
Define:
$\Delta_{\text{main}}=\mathcal{L}_{\text{main}}(\theta;x,y)-\mathcal{L}_{\text{main}}(\theta^{\prime};x,y),$
the difference in main task loss due to test-time updates:
$\Delta_{\text{main}}\approx\nabla_{\theta}\mathcal{L}_{\text{main}}\cdot\Delta\theta,$
where $\Delta\theta=-\eta\nabla_{\theta}\mathcal{L}_{\text{self}}(\theta;x)$.
Since $\mathcal{L}_{\text{self}}$ is orthogonal to $\mathcal{L}_{\text{main}}$ by design,
$\nabla_{\theta}\mathcal{L}_{\text{main}}\cdot\nabla_{\theta}\mathcal{L}_{\text{self}}\approx 0,$
leading to negligible interference with the main task.
The claim that $\mathcal{L}_{\text{self}}$ (the self-supervised task loss) is orthogonal to $\mathcal{L}_{\text{main}}$ (the main task loss) is a simplifying assumption that holds in certain cases due to how the self-supervised tasks are typically designed. Below we present the reasoning behind this assumption and its justification.
Why Orthogonality is Assumed
  1. 1.
    Distinct Optimization Objectives:
    • Self-supervised tasks ($\mathcal{L}_{\text{self}}$) are often designed to exploit auxiliary structures or representations in the data (e.g., rotation prediction, reconstruction).
    • Main tasks ($\mathcal{L}_{\text{main}}$) are task-specific and rely on labeled data.
    • By design, $\mathcal{L}_{\text{self}}$ operates on a different objective that does not directly interfere with $\mathcal{L}_{\text{main}}$.
  2. 2.
    Gradient Independence:
    • The gradients $\nabla_{\theta}\mathcal{L}_{\text{self}}$ and $\nabla_{\theta}\mathcal{L}_{\text{main}}$ are computed from different aspects of the model’s output.
    • For example, if $\mathcal{L}_{\text{self}}$ reconstructs data and $\mathcal{L}_{\text{main}}$ classifies labels, their parameter updates are unlikely to point in similar directions.
Formalization of Orthogonality
The assumption of orthogonality can be expressed as:
$\nabla_{\theta}\mathcal{L}_{\text{main}}\cdot\nabla_{\theta}\mathcal{L}_{\text{self}}\approx 0.$
This condition implies that:
$\cos\theta\approx 0,$
where $\theta$ is the angle between the gradient vectors.
Justification for Approximate Orthogonality
  1. 1.
    Design Choice: Self-supervised tasks are chosen to be auxiliary and independent from the main task. For instance:
    • Rotation Prediction (self-supervised) vs. Classification (main task): Gradients act on different representations.
    • Reconstruction Tasks: Focus on encoding input features rather than task-specific labels.
  2. 2.
    Empirical Evidence: In [33], the authors show that TTT optimizations during inference generally improve robustness without significantly altering the main task’s performance. This is indirect evidence that the gradient interference is minimal.
  3. 3.
    Gradient Magnitudes: Test-time updates often involve small gradient steps ($\eta\ll 1$), making any interference negligible.
When Orthogonality Might Not Hold
  • If the self-supervised task is too closely related to the main task, gradient overlap can occur, leading to interference.
  • If the auxiliary task introduces biases that affect the features used by the main task, orthogonality breaks down.
Numerical Verification
To empirically check for orthogonality:
  1. 1.
    Dot Product Test: Compute the dot product of the gradients:
[…]
    If the result is close to zero, the tasks are approximately orthogonal.
  2. 2.
    Loss Curve Analysis: Monitor the changes in $\mathcal{L}_{\text{main}}$ during self-supervised updates:
    $\Delta\mathcal{L}_{\text{main}}=\mathcal{L}_{\text{main}}(\theta)-\mathcal{L}_{\text{main}}(\theta^{\prime}),$
    where $\theta^{\prime}$ is updated using $\mathcal{L}_{\text{self}}$. Minimal changes imply negligible interference.
A.3 Handling Extreme Distribution Shifts and Computational Overhead
TTT leverages self-supervised tasks invariant to distribution shifts, such as rotation prediction or reconstruction tasks. These tasks guide the model to reorient itself in a new feature space without requiring explicit labels.
Mathematical Adaptation Framework:
Under extreme distribution shifts, let $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ denote the training and test distributions, respectively, such that:
$\mathcal{D}_{\text{train}}\neq\mathcal{D}_{\text{test}}.$
The goal is to adapt the model $f_{\theta}$ to the shifted distribution $\mathcal{D}_{\text{test}}$ using:
$\mathcal{L}_{\text{self}}(\theta;\mathbf{x})=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}_{\text{test}}}\left[\text{auxiliary loss}(\mathbf{x})\right].$
Proof of Robustness to Distribution Shifts:
Let $P_{\text{train}}(\mathbf{x})$ and $P_{\text{test}}(\mathbf{x})$ represent the training and test data distributions. Using auxiliary tasks, TTT minimizes:
[…]
The minimization of $\mathcal{L}_{\text{self}}$ aligns $f_{\theta}$ with $P_{\text{test}}$, adapting the model to the test distribution.
Computational Overhead:
For each test sample $\mathbf{x}$, the overhead is:
  1. 1.
    Forward pass on $\mathcal{L}_{\text{self}}$: $\mathcal{O}(T\cdot d)$.
  2. 2.
    Backpropagation to compute gradients: $\mathcal{O}(U\cdot T\cdot d^{2})$.
Total per-sample overhead: $\mathcal{O}(U\cdot T\cdot d^{2})$.
A.4 Impact on Computational Resources
Memory Usage:
Let $M_{\text{model}}$ denote the base memory required for the model:
  • Test-time gradients increase memory usage proportional to $T\cdot d$:
    $M_{\text{TTT}}=M_{\text{model}}+\mathcal{O}(T\cdot d).$
Latency and Runtime:
Test-time updates introduce additional runtime:
$\text{Latency}_{\text{TTT}}=\text{Latency}_{\text{model}}+\mathcal{O}(U\cdot T\cdot d^{2}),$
where $U$ is the number of iterations for test-time optimization.
Proof of Impact:
Define the test-time computation for $\mathcal{L}_{\text{self}}$ as:
$C_{\text{TTT}}=\text{Forward}_{\text{self}}+\text{Backward}_{\text{self}}.$
  • Forward: $\mathcal{O}(T\cdot d)$ (invariant to the base model complexity).
  • Backward: $\mathcal{O}(U\cdot T\cdot d^{2})$.
A.5 Parameter Initialization in TTT
Test-Time Training (TTT) does not require specialized or customized parameter initialization methods. For backbone architectures, TTT modules utilize standard initialization techniques, such as Xavier or He initialization, to ensure stable learning dynamics. Since TTT’s test-time updates are based on the weights learned during training, the model is agnostic to specific initialization strategies.
While TTT does not mandate particular initialization methods, it can benefit from pretrained weights. By using a pretrained backbone, the model can leverage representations already optimized for a related domain, allowing the test-time updates to refine these representations further. For example, substituting a pretrained backbone with a TTT module can enhance adaptability during inference without requiring substantial retraining.
Empirical results from prior studies (e.g., [33]; [32]) support this observation. While pretrained weights can enhance performance, they are not strictly necessary. TTT’s adaptability and effectiveness primarily stem from its self-supervised task, which guides the model to align with the test distribution rather than relying on the initialization strategy.
This demonstrates that TTT is flexible and performs well across different initialization settings, with its core strength being its adaptability at test time. Further elaboration on this topic can be found in the cited references.
A.6 Generalization of TTT Beyond Time Series Forecasting
Furthermore, we wish to emphasize that TTT generalizes well beyond time series forecasting. From [32], TTT has been successfully applied to Language Modeling, where it demonstrated competitive results compared to Mamba and Transformer-based models. In [33], TTT was applied to Object Recognition, where it improved performance on noisy and accented images in the CIFAR-10-C dataset by adapting at test time. Finally, in [35], TTT was extended to Video Prediction, enabling the model to adjust to environmental shifts such as changes in lighting or weather.
These works collectively illustrate the generalization of TTT to other sequence modeling tasks and its effectiveness across diverse domains, including Vision Prediction, Language Modeling, and Object Recognition apart from Time Series Forecasting.
A.7 Failure Case Study
In [33], TTT was tested on CIFAR-10-C, a corrupted version of CIFAR-10 that includes 15 types of distortions such as Gaussian noise, motion blur, fog, and pixelation applied at five severity levels. These corruptions simulate significant distribution shifts from the original dataset. The results demonstrated that TTT significantly improved classification accuracy, achieving 74.1% accuracy compared to 67.1% accuracy for models that did not adapt during test time.
Notably:
  • Under severe shifts like Gaussian Noise, TTT effectively adapted to noisy inputs, outperforming baseline models that lacked test-time updates.
  • For distortions like motion blur and pixelation, TTT successfully reoriented the model’s feature space to handle spatial distortions.
Compared to methods such as domain adaptation and augmentation-based approaches, TTT demonstrated superior performance under extreme distribution shifts, highlighting its robustness and adaptability.
While these results focus on image classification, they provide strong evidence of TTT’s capability to handle abrupt distributional changes, which can be analogous to sudden anomalies in time series data. We acknowledge that a failure case analysis specific to Time Series Forecasting is a valuable avenue for future research and appreciate the reviewer’s suggestion.
For more detailed results, we encourage the reader to refer to [33].
Appendix B TTT vs Mamba
@@
B.1 Mamba: Gated Linear RNN via State Space Models (SSMs)
@@
B.2 TTT: Test-Time Training with Dynamic Hidden States
@@
Comparison of Complexity and Adaptability
@@
Appendix C Comparisons with Models Handling Noise or Temporal Regularization
C.1 Comparisons with Models Handling Noise or Temporal Regularization
To position Test-Time Training (TTT) relative to the state-of-the-art, we compare its performance with models specifically designed for noise robustness or temporal regularization:
Comparison with DeepAR
DeepAR is a probabilistic forecasting model that handles uncertainty in time series data using autoregressive distributions. While it excels in forecasting under stochastic conditions, TTT’s test-time adaptation offers significant advantages in handling sudden, unseen distributional shifts.
Comparison with TCN (Temporal Convolutional Network)
Temporal Convolutional Networks (TCNs) are known for their ability to capture long-range dependencies efficiently. However, TCNs lack the adaptability of TTT, which dynamically aligns feature representations during test time. Adding noise to datasets like ETTh1 or ETTm1 could further highlight TTT’s advantage over static methods such as TCN.
C.2 Theoretical Comparison
1. Static Models (e.g., DeepAR, TCN)
Static models like DeepAR and TCN rely on fixed parameters $\theta$ that are learned during training and remain unchanged during inference. Mathematically:
$\hat{\mathbf{y}}=f_{\theta}(\mathbf{x}),$
where $\mathbf{x}$ represents the input sequence, $\hat{\mathbf{y}}$ is the forecasted output, and $f_{\theta}$ is the model with fixed parameters $\theta$.
These models excel under stationary conditions or when the training and testing distributions $P_{\text{train}}(\mathbf{x})$ and $P_{\text{test}}(\mathbf{x})$ are similar. However, they struggle under distribution shifts, where $P_{\text{train}}(\mathbf{x})\neq P_{\text{test}}(\mathbf{x})$, as they cannot adapt their parameters to align with the shifted test distribution.
2. TTT’s Dynamic Adaptation
Test-Time Training (TTT) introduces a test-time adaptation mechanism that updates the model parameters dynamically based on a self-supervised loss. During inference, the parameters $\theta$ are updated as follows:
$\theta^{\prime}=\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{self}}(\theta;\mathbf{x}),$
where:
  • $\mathcal{L}_{\text{self}}(\theta;\mathbf{x})$ is the self-supervised auxiliary loss designed to align the model’s representations with the test distribution.
  • $\eta>0$ is the learning rate for test-time updates.
This dynamic adjustment allows TTT to adapt to unseen distribution shifts $P_{\text{test}}(\mathbf{x})$ by optimizing the feature representations for each test sample, resulting in improved generalization:
$\hat{\mathbf{y}}=f_{\theta^{\prime}}(\mathbf{x}),$
where $\theta^{\prime}$ is dynamically updated for each test instance. This mechanism enables TTT to handle abrupt, non-stationary shifts that static models cannot address effectively.
3. Comparison of Noise Robustness
To further compare, consider a scenario with noisy inputs $\mathbf{x}+\epsilon$, where $\epsilon\sim\mathcal{N}(0,\sigma^{2})$.
Static Models: The forecast relies on fixed parameters:
$\hat{\mathbf{y}}_{\text{static}}=f_{\theta}(\mathbf{x}+\epsilon).$
Without adaptive mechanisms, noise $\epsilon$ directly degrades the model’s performance, as the learned parameters $\theta$ are not optimized for the noisy distribution.
TTT: TTT updates its parameters to account for the noisy inputs:
$\theta^{\prime}=\theta-\eta\nabla_{\theta}\mathcal{L}_{\text{self}}(\theta;\mathbf{x}+\epsilon).$
This update minimizes the impact of $\epsilon$ by dynamically realigning the feature representations, resulting in improved predictions:
$\hat{\mathbf{y}}_{\text{TTT}}=f_{\theta^{\prime}}(\mathbf{x}+\epsilon).$
Empirically, this adaptability enables TTT to outperform static models in scenarios with noise or abrupt distribution shifts.
4. Summary
The key difference lies in the adaptability:
  • Static models like DeepAR and TCN rely on fixed parameters and are effective under stationary conditions but struggle with non-stationary data or noise.
  • TTT dynamically adjusts its parameters using self-supervised learning at test time, providing a significant advantage in handling distribution shifts and noisy inputs.
Appendix D Model Components
D.1 TTT Block and Proposed Architectures
@@ D.3 Normalization
D.4 Expanding on the Choice of Hierarchical Two-Level Context Modeling
The hierarchical design of Test-Time Training (TTT) is well-suited for tasks like time series forecasting due to its ability to adapt across both high- and low-resolution contexts. Below, we outline the benefits of this structure for time-series forecasting:
Hierarchical Representation of Temporal Dependencies
Multiscale patterns in time series data, such as daily, weekly, or seasonal trends, require capturing both fine-grained and coarse-grained temporal dependencies. Architectures like Conv Stacked 5 and ModernTCN implicitly model hierarchical temporal features through stacked convolutional layers and depthwise-separable convolutions, respectively. These architectures balance local temporal feature extraction with the global adaptability provided by TTT.
Adaptation to Non-Stationary Patterns
The hierarchical design ensures that the model can adapt to distribution shifts occurring at different temporal resolutions. For example:
  • Sudden anomalies in fine-grained data.
  • Gradual trends in coarse-grained data.
Proposed Benchmarks
To validate TTT’s ability to adapt to multiscale patterns, we propose the following:
  • Additional evaluations on noise-robust datasets, such as adding noise to ETTh1 and ETTm1.
  • Temporal regularization tasks using benchmarks like DeepAR or Prophet, which can serve as strong baselines for comparison.
Appendix E Analysis on Increased Prediction and Sequence Length
E.1 Effect of Sequence Length